### v2

# Linguistics and Human Brain: A Perspective of Computational Neuroscience

Abstract
Elucidating the language–brain relationship requires bridging the methodological gap between linguistics’ abstract theoretical frameworks and neuroscience’s empirical neural data. As an interdisciplinary cornerstone, computational neuroscience formalizes language’s hierarchical and dynamic structures into testable neural representation models through modeling, simulation, and data analysis, enabling computational dialogue between linguistic hypotheses and neural mechanisms. Recent advances in deep learning, particularly large language models (LLMs), have further advanced this inquiry: their high-dimensional representational spaces provide a new scale for probing the neural basis of linguistic processing, while the model–brain alignment framework offers a principled approach to evaluating the biological plausibility of language-related theories.
This review synthesizes interdisciplinary progress from a computational neuroscience perspective. First, it outlines the core connotations of major linguistic frameworks (generative grammar, functional linguistics, and cognitive linguistics), their cross-cultural and evolutionary characteristics, and key challenges for neural alignment, including limited quantitative mechanisms, poor accessibility of abstract constructs to neural measures, and insufficient treatment of dynamics and plasticity. Second, it introduces the methodological foundations of linguistics–neuroscience dialogue, focusing on four technical pillars: neural activity measurement (e.g., fMRI, EEG, MEG, fNIRS, ECoG, SEEG), linguistic numerical representation, the evolution of language models from statistical approaches to LLMs, and neural coding frameworks that link model representations to brain signals, illustrated with a model–brain alignment case study. Third, it summarizes major findings, ranging from early computational insights into predictability and structural processing to recent LLM-driven progress in cross-modal interaction, inter-brain coupling, hierarchical computation, learning strategy sensitivity, and language plasticity. Finally, the review discusses current limitations—including functional alignment without structural homology, constraints on real-time validation, biased research coverage, and narrow evaluation metrics—and proposes future directions, such as improving biological plausibility via spiking neural network–based language models, developing cognitive-level alignment frameworks integrating memory, causality, and metacognition, and extending clinical applications.
In summary, this work aims to advance a comprehensive, mechanistic understanding of the language-brain relationship and promote computational neuroscience as a generative theoretical framework for testable neuro-computational accounts of language.
Keywords Linguistics $\cdot$ Brain $\cdot$ Computational Neuroscience $\cdot$ Neuroimaging $\cdot$ Neural Coding $\cdot$ Brain-Computer Interface $\cdot$ Language Model
1 Introduction
Language is conceptualized as a multi-layered abstract symbolic system encompassing distinct yet interconnected structures ranging from sounds (phonetics and phonology) to word formation (morphology), sentence structure (syntax), and meaning (semantics) [35, 34]. Using a finite set of discrete elements and combinatorial rules, it can generate an infinite array of expressions, enabling the flexible and precise transmission of meaning – a feature known as recursion [20, 72, 94, 37]. Linguistics seeks to formalize these implicit rules and structures in order to uncover the cognitive architecture underlying human linguistic competence. In parallel, neuroscience investigates how coordinated neural activity across brain circuits implements the processes of language production, processing and comprehension [107, 216]. A persistent interdisciplinary challenge, however, arises from the methodological and explanatory divide between these fields. Abstract linguistic theories are often formulated as symbolic, hierarchical systems that are difficult to map directly onto the dynamic, distributed patterns of neural activity observed through neuroimaging or electrophysiology, partly due to the inherent limitations of these techniques. Conversely, neural data alone often lack the computational interpretability needed to account for the structured, rule-governed nature of language. This gap between theoretical description and empirical evidence limits a comprehensive understanding of language and its neural basis.
Beyond methodological mismatches, a more fundamental difficulty lies in the nature of the neural system itself. The neural system for language is a complex, adaptive, and dynamic network, composed of billions of neurons that exhibit intricate connectivity and continuously evolving plasticity [76, 38]. Within this network, linguistic information is processed through hierarchical, parallel, and recurrent interactions, supported by bidirectional inter-regional connections that are dynamically regulated by contexts and task demands [55, 170]. In formal terms from mathematics and systems science, such a system is described as an adaptive complex dynamic system [133] where global behavior cannot be reduced to the sum of its local parts, and it often exhibits nonlinear, self-organizing, and emergent properties. Consequently, neither purely linguistic models nor isolated neural observations can fully explain the integrated mechanisms of human language processing.
As a well-founded discipline dedicated to decoding neural mechanisms, computational neuroscience serves as a crucial bridge between linguistics and neuroscience. It integrates linguistics, neuroscience, computer science, and systems theory to convert formal linguistic hypotheses into testable computational models, which are then tested against neural data [104, 87]. The core methodology involves constructing models that represent linguistic structure while simulating neural dynamics, which are then validated or revised using brain imaging and electrophysiological evidence [147, 124]. Ever since generative linguistics posited a neural basis for linguistic competence [136], computational modeling has become a central tool for evaluating the neural plausibility of linguistic theories.
Recent advances in artificial intelligence, particularly in deep learning and Large Language Models (LLMs), have accelerated this integration further. Researchers can now explore linguistic structures and their correspondence with brain networks within higher-dimensional representational spaces [226, 214, 247]. LLMs provide computational platforms for examining modern linguistic phenomena such as semantic integration, long-distance dependencies, and predictive processing. They also support the emerging framework of model–brain alignment, wherein internal model representations are used to predict and explain neural responses during language processing [56, 5, 117]. This development marks a shift from early conceptual modeling toward quantitative mappings between large-scale linguistic data and high-resolution neural signals.
This paper provides a systematic review of how computational neuroscience serves as a methodological bridge between linguistics and neuroscience. Section 2 surveys modern linguistic frameworks and their development across languages, identifying the core challenges that have historically hindered their alignment with neural data. Section 3 introduces key methodological foundations, including neural data acquisition, word embeddings, modern language models, and neural coding frameworks. Section 4 reviews applications of classical computational neuroscience models to language comprehension, followed by a synthesis of recent progress in LLM-driven neural alignment, covering domains including cross-modal representation, cross-brain coupling, hierarchical computation, and learning mechanisms. Section 5 analyzes current methodological limitations and theoretical challenges, and outlines promising future research directions, such as enhancing biological plausibility, developing cognitive-level alignment frameworks, and expanding clinical and brain–computer interface applications. Finally, Section 6 summarizes the main conclusions and discusses the long-term significance of computational neuroscience for fostering a deeper integration of linguistics and neuroscience.
2 Theoretical Foundations of Linguistics
Language has long been the core research object defining the intersection of linguistics and neuroscience. Over time, linguistic theories have continued to develop through modeling, cross-cultural research, and empirical investigation. Yet they have also gradually revealed structural limitations that hinder their direct alignment with the neural mechanisms of the brain. Modern linguistic theories offer clear, organized ways to describe the structure and function of language. But when faced with the diversity of languages across cultures, the dynamic nature of everyday language use, and growing evidence from brain science, their ability to fully explain language remains limited [70, 60, 162, 85]. In this section, we review major linguistic perspectives and their evolution across cultural contexts, identifying the key challenges that motivate interdisciplinary integration. This review provides the theoretical background for later discussions on computational neuroscience as a bridging framework, as outlined in Fig. 1.
Figure 1: Modern Linguistic Theories and Challenges for Bridging Linguistics and Neuroscience.
2.1 Frameworks of Modern Linguistic Theories
Modern linguistics comprises several influential theoretical frameworks. Among the most influential are generative grammar, functional linguistics, and cognitive linguistics, each offering a distinct perspective on language structure and use.
Generative grammar, most prominently associated with Noam Chomsky, posits linguistic competence as an innate, biologically grounded component of the human cognitive system [35]. This framework is founded on the hypothesis of a Universal Grammar, arguing that children’s acquisition of complex grammar must be guided by innate principles, given the impoverished and limited nature of their linguistic input. This "poverty of the stimulus" argument is taken as primary evidence for internally constrained, domain-specific grammatical principles. The distinction between deep structure (abstract syntactic relations) and surface structure (observable linguistic forms) illustrates how underlying representations are transformed into observable linguistic expressions [36, 109]. Later developments, most notably the Minimalist Program, sought to reduce linguistic theory to a minimal set of core computational operations, with Merge as the fundamental process, to enhance its cognitive and biological plausibility [134]. Nevertheless, these highly abstract, formal constructs have proven difficult to map onto specific, observable patterns of neural activity, which remains a source of sustained debate in the neurobiology of language.
In contrast, functional linguistics treats language primarily as a system for social communication. Its central tenet is that linguistic structure is shaped and constrained by communicative needs and contextual factors [237]. A major framework within this tradition, Systemic Functional Linguistics, posits that language simultaneously fulfills three core metafunctions: the ideational (for construing experience), the interpersonal (for enacting social relations), and the textual (for organizing discourse). Within this framework, linguistic choices are fundamentally motivated by a speaker’s communicative goals and the dynamics of social interaction [89]. Structural differences between language modalities, such as the prevalence of ellipsis in spoken conversation versus the structural completeness expected in formal writing, are analyzed as adaptive outcomes of the distinct functional demands of each modality [65]. While functional linguistics has been highly influential in discourse analysis, sociolinguistics, and cross-cultural communication studies, its strong emphasis on usage, context, and meaning poses significant challenges for strict formalization. This inherent difficulty with formal modeling, in turn, complicates the establishment of direct, mechanistic links between functional explanations and underlying neural processes.
Cognitive linguistics explicitly rejects the generative conception of language as an autonomous, innate cognitive module. Instead, it argues that language is grounded in, and emerges from, general cognitive processes such as categorization, metaphorical mapping, and mental simulation [130]. Within this framework, Conceptual Metaphor Theory proposes that abstract domains (e.g., time, mind) are understood through systematic mappings from more concrete, embodied experiences, and that grammatical structures often reflect these underlying conceptual representations [122]. Embodied cognition accounts extend this view, proposing that language comprehension is inherently tied to the reactivation of sensorimotor experiences and dynamic mental simulation [9]. While this perspective offers a powerfully unified account of language and cognition, many of its core constructs, such as conceptual metaphors or mental simulations, are difficult to operationalize and quantify experimentally. This lack of form specification consequently limits the framework’s ability to generate precise, testable predictions that can be directly compared with neural data.
Overall, these three major theoretical frameworks account for language in terms of innate mechanisms, communicative function, and general cognition, respectively. However, their explanations remain largely abstract. A substantial gap persists between high-level linguistic descriptions and the concrete neural implementation of language. This gap poses a major challenge for both computational modeling and neurobiological investigation.
2.2 Cross-Cultural Evolution of Linguistic Theories
The existence of thousands of languages worldwide exhibits remarkable diversity in linguistic structure and expression. This diversity has continuously driven the revision and expansion of linguistic theory, while also underscoring the challenge of constructing a unified explanatory framework, raising questions about whether genuinely universal linguistic features can be abstracted at all [171, 47].
Early generative grammar was primarily developed on the basis of a small number of languages. To account for typological variation, it later introduced parameter-setting mechanisms to explain differences in word order, morphological structure, and subject realization [34]. However, for languages with highly complex structures or those that differ substantially from Indo-European languages, parameter-based explanations remain controversial. This suggests that a limited set of parameters may be insufficient to capture the full richness of linguistic diversity [171].
In functional linguistics, cross-linguistic variation is examined in relation to the communicative and cultural contexts in which language is used. Honorific systems, context-dependent expressions, and discourse-level differences are treated not as peripheral but as core objects of grammatical analysis [89]. For example, cultures differ in their emphasis on politeness, indirectness, and social hierarchy, which in turn shape linguistic forms [10]. Although this perspective helps explain cultural variation, its strong reliance on specific cultural contexts makes it difficult to extract universally applicable principles.
Cognitive linguistics examines the interaction between universal embodied experience and culturally specific conceptualizations [130]. Many spatial and emotional metaphors are shared across languages (e.g., time is conceived along a front-back axis, and emotion is metaphorically mapped onto a vertical axis, with happiness represented as "up" and sadness as "down"), suggesting shared cognitive constraints. At the same time, domains such as spatial reference (with systems based on body coordinates like left/right/front/back or on cardinal directions like north/south/east/west), color categorization (with the number and boundaries of basic colour terms varying across languages), and kinship terminology (with varying distinctions between nuclear and extended family) exhibit clear cultural specificity [121]. The coexistence of universality and diversity indicates that language is shaped by both biological constraints and cultural experience.
Cross-linguistic research thus demonstrates that language structure reflects shared cognitive foundations as well as cultural history [39]. However, this twofold origin poses a challenge for determining which linguistic properties stem from universal neural mechanisms and which are products of cultural transmission. Evidence from cross-script studies of developmental dyslexia further illustrates this complexity: reading impairments in alphabetic versus logographic systems engage partially distinct neural circuits, suggesting that the neural correlates of dyslexia reflect both universal constraints and writing-system-specific adaptations [220]. Such nuanced findings underscore the need for closer integration between theoretical linguistics and neurobiologically grounded models of language.
2.3 Challenges in Modern Linguistic Theories
With advances in neuroimaging techniques and computational methods, language research has increasingly moved beyond purely theoretical analysis toward empirical integration with neural data [87]. In this interdisciplinary context, major linguistic theories have revealed several structural challenges. These challenges limit direct links between linguistic constructs and neural mechanisms while motivating the development of new computational and neural modeling approaches [238]. Broadly speaking, these difficulties include a lack of quantitative mechanisms, limited research scope, and poor integrability of findings.
  1. i.
    Lack of explicit quantitative mechanisms
    Most linguistic and cognitive language studies remain at the level of phenomenological description and qualitative explanation. The computational mechanisms underlying language processing are often underspecified. For example, during lexical processing, the N400 event-related potential (ERP) is frequently observed, and numerous studies have shown that its amplitude is modulated by contextual information [51, 25]. However, such findings typically demonstrate the existence of contextual effects without clarifying the underlying computations.
    Several interpretations have been proposed in the cognitive science literature. Some link the N400 to lexical predictability, suggesting that more predictable words elicit smaller responses. Others interpret it as reflecting semantic integration difficulty [157]. Until recently, these accounts remained largely conceptual, lacking specification of how prediction and integration are implemented computationally within neural circuits. While recent computational models have begun to address this gap, no unifying theory has yet achieved consensus, and the challenge of precise theory-data correspondence remains central to the field.
  2. ii.
    Research restricted to specific linguistic phenomena
    Language research often relies on highly controlled experimental designs to ensure reliability and interpretability [240]. Such designs typically focus on decontextualized or localized linguistic phenomena using tightly constrained, uniform stimuli. As a result, findings may not generalize to the richness and variability of natural language use.
    A further challenge lies in the long-standing misalignment of research scales between linguistics and neuroscience [49]. Linguistic studies often examine fine-grained structural distinctions, whereas neuroscience research typically focuses on broader functional organization across brain regions. Even in interdisciplinary work, experiments predominantly rely on simplified or prototypical materials, limiting ecological validity and generalizability.
  3. iii.
    Difficulty in forming a unified explanatory framework
    While controlled experiments are effective at identifying local effects, they also contribute to fragmented findings. Individual studies typically target specific phenomena, yet language itself resists decomposition into a small set of independent units. Differences in task design, stimulus materials, and measurement techniques further hinder direct integration across studies. For instance, distinct types of linguistic stimuli may elicit divergent activation patterns or electrophysiological responses. However, aggregating such results rarely yields a coherent account of how the brain incrementally constructs the meaning of a complete sentence [99].
  4. iv.
    Abstract linguistic constructs are difficult to map onto neural measures
    Many central linguistic concepts, such as recursion, thematic roles, and conceptual metaphor, are highly abstract and lack clear operational definitions, making them difficult to translate into measurable neural variables [193]. Moreover, neural activity is distributed across large-scale networks rather than confined to single regions, further complicating attempts to establish simple, one-to-one mappings between linguistic functions and specific neural indicators [62].
  5. v.
    Lack of a unified explanation for cross-linguistic differences
    There is no consensus on whether cross-linguistic structural differences arise primarily from innate mechanisms or from cultural learning and experience. These divergent theoretical positions make it difficult for neuroimaging studies to determine whether observed processing differences reflect biological predispositions or experiential adaptation. This uncertainty hinders the development of a unified neurocognitive model of language.
  6. vi.
    Limited ability to explain dynamics and plasticity in language processing
    Major linguistic theories are largely based on static sentence analysis and struggle to account for the dynamic aspects of real-world communication, including real-time prediction, contextual updating, and social interaction processes under naturalistic conditions [189, 188]. In addition, human language abilities show substantial plasticity during child acquisition, second language learning, and recovery after brain injury. These theoretical frameworks lack a unified account of these developmental and adaptive changes, as well as of individual differences in language processing.
Taken together, abstract linguistic theories alone are insufficient for establishing direct links between language and neural mechanisms. Bridging this gap requires translating linguistic constructs into computable models and testing them quantitatively against neural data. Computational neuroscience has emerged in this context as a key platform for connecting linguistic theory with brain mechanisms. By integrating neural data, computational modeling, and large-scale models of language processing, this field is promoting a shift in language research from descriptive accounts toward mechanistic explanations and laying a critical foundation for subsequent work on brain-inspired models and advances in language processing.
Figure 2: The Association between Linguistics and Neuroscience from the Perspective of Computational Neuroscience.
3 Computational Neuroscience Methods for Linguistics
Computational neuroscience provides a methodological bridge between formal linguistic theory and empirical neural data. As illustrated in Fig. 2, it transforms abstract linguistic structures into computational representations that can be quantitatively compared with neural activity through a set of interrelated tools and analytical frameworks. This section introduces the core components of this bridge. We first outline the methodology of computational neuroscience and then describe four technical pillars: Measurement Methods (neural activity recording), Word Embeddings (numerical representations of language), Evolving Language Models, and Neural Coding (analytical frameworks that link model representations to brain signals). Finally, we present an integrative case study to illustrate how these components operate jointly and support model–brain alignment research in linguistics and neuroscience.
3.1 Methodology in Computational Neuroscience
As an interdisciplinary field, computational neuroscience adopts diverse and integrative methodological strategies. Rather than pursuing a single unifying theory, it typically develops modular and composable models to study complex systems such as the language and brain [95]. Model evaluation emphasizes explanatory and predictive power with respect to neural data, formal parsimony, and the ability to generate testable hypotheses [139]. From this perspective, modeling approaches to language research have traditionally been grouped into three categories:
  • Descriptive models, which quantify observed neural activity, including spike train statistics and population decoding methods [52, 204, 176, 194, 73].
  • Normative theories, which explain neural computation through functional optimization principles, such as efficient coding and Bayesian inference [7, 217, 116, 119, 97].
  • Biological simulation models, which aim to reproduce neural structure and dynamics at varying levels of biological detail [153, 66, 22].
The emergence of deep learning models, particularly LLMs, challenges this traditional classification. LLMs are not designed to fit neural data or obey biological constraints. Nevertheless, their strong correspondence with human language behavior, representational structure, and neural responses positions them as a distinctive mesoscale computational reference frame [124, 212, 4, 82, 30, 246, 28, 29, 199, 177]. Their central contribution lies in offering an explicit, hierarchical representational space that enables systematic investigation of how linguistic information may be organized in the brain [61].
These developments motivate a new framework of multi-scale integration. Within this framework, LLMs characterize linguistic structure at the computational level, cognitive interpretability is introduced at the algorithmic level, and biological constraints are progressively incorporated at the implementation level. Accordingly, the primary research goal shifts from fitting neural data toward establishing principled connections between expressive computational models and neurobiological mechanisms. This shift supports progress from black-box modeling toward mechanistic explanation.
Table 1: Core components and their roles in model-brain alignment research.
Core Components Core Forms/Technologies Core Functions Roles in Alignment
Measurement Methods fMRI, EEG, MEG, ERP Recording neural activity signals Providing neural reference data
Word Embedding High-dimensional semantic vector mapping Numerical representation of linguistic units Laying the foundation for linguistic representation
Evolving Language Models n-gram, Transformer architecture, etc. Learning intrinsic linguistic rules Providing computational reference frame
Neural Coding Regression/Neural network models Establishing representation-neural mapping Quantifying model-brain matching degree
Guided by this methodological perspective, the following sections introduce the four technical pillars of this framework in detail and clarify their respective roles in model–brain alignment. Their respective roles and interactions in model–brain alignment research are summarized in Table 1.
3.2 Brain Activity Measurement for Linguistic Tasks
Testing linguistic theories within a computational framework requires reliable measurements of neural activity during language processing. Research on the neural basis of language therefore employs multiple techniques that capture neural responses across distinct spatial, temporal, and representational scales [99, 75, 236, 113, 84]. Because language processing emerges from dynamic interactions across distributed brain regions and multiple scales, no single method is sufficient. Instead, measurement techniques are selected and combined according to the research focus, such as spatial localization, temporal resolution, or representational specificity.
Table 2: Comparison of characteristics and applicable scenarios of mainstream techniques for detecting neural mechanisms of language.
Measurement Techniques Spatial Resolution Temporal Resolution Core Advantages Applicable Scenarios in Language Research Limitations
fMRI Millimeter-level Second-level Precise brain region localization Brain region distribution at linguistic levels Unable to capture millisecond-scale dynamics
EEG Centimeter-level Millisecond-level Capturing temporal dynamics Instantaneous responses to semantics/syntax Ambiguous spatial localization
MEG Sub-centimeter-level Millisecond-level Balanced spatiotemporal resolution Mechanisms of phonetic-lexical conversion High equipment cost
ERP Centimeter-level Millisecond-level Extraction of event-related components Responses to specific linguistic phenomena Dependent on multiple stimulus superposition
fNIRS Centimeter-level Hundred-millisecond-level Strong anti-motion artifact performance Natural scene language communication Limited penetration depth
ECoG Millimeter-level Sub-millisecond-level High SNR, fine cortical spatial precision Presurgical language mapping, cortical language dynamics Invasive, only for clinical patients
SEEG Millimeter-level Sub-millisecond-level Deep structure 3D localization, high precision Subcortical-cortical language networks, deep language circuits Invasive, surgical risk, limited coverage
In this section, we introduce seven widely used brain activity measurement techniques in language research. Their core characteristics and methodological trade-offs are summarized in Table 2.
  1. i.
    fMRI (functional Magnetic Resonance Imaging): fMRI is a core non-invasive technique for measuring brain function based on the Blood Oxygen Level–Dependent (BOLD) effect [255, 174]. Neural activation increases metabolic oxygen consumption, which triggers compensatory increases in local cerebral blood flow. The resulting influx of oxygenated blood exceeds immediate neuronal demand, reducing local deoxyhemoglobin concentration. Because deoxyhemoglobin is paramagnetic whereas oxyhemoglobin is diamagnetic, these changes modify local magnetic resonance signals, allowing indirect imaging of regional neural activation.
    The main advantage of fMRI is its high spatial resolution, which enables accurate localization of cortical and subcortical regions involved in language processing, including left perisylvian areas (Broca’s area, Wernicke’s area, supramarginal gyrus, angular gyrus) and subcortical structures such as the caudate nucleus and thalamus. It can also distinguish spatial activation patterns associated with different linguistic levels such as phonological, lexical, semantic, and syntactic processing [18, 195, 19, 241, 256]. Its primary limitation is low temporal resolution: the BOLD response is delayed by approximately 4–8 seconds, with signal peaks typically occurring 6–8 seconds after neuronal activation. Consequently, millisecond-scale processes such as rapid semantic integration cannot be directly captured [80, 154]. In addition, fMRI measurements are sensitive to physiological noise from cardiac and respiratory activity, which complicates experiments involving infants or patients with language disorders. A further limitation is that the BOLD signal is an indirect correlate of neural activity, reflecting hemodynamic changes rather than neuronal firing itself. Overall, fMRI offers high spatial resolution but limited temporal resolution [16].
  2. ii.
    EEG (Electroencephalography): EEG records weak electric field fluctuations generated by the synchronous firing of large neuronal populations using scalp electrode arrays [140]. Its key advantage is millisecond-level temporal resolution, which enables precise tracking of rapid neural dynamics during language processing, including responses related to semantic and syntactic analysis. However, electrical signals are attenuated and spatially blurred when the pass through brain tissue, skull, and scalp, resulting in poor spatial localization and significant difficulty in identifying precise neural sources [172, 152]. EEG is therefore highly sensitive to synchronized cortical activity but provides limited spatial precision [149]. This positions EEG as a complement to fMRI, which offers high spatial resolution but limited temporal resolution.
  3. iii.
    MEG (Magnetoencephalography): MEG measures weak magnetic fields generated by neuronal electrical activity using highly sensitive sensors – traditionally Superconducting Quantum Interference Devices (SQUIDs) [44, 90] and, more recently, Optically Pumped Magnetometers (OPMs). Synchronous neuronal currents produce magnetic fields that can be detected outside the head, providing an indirect measure of neural activity. Like EEG, MEG offers millisecond temporal resolution and can track rapid processing stages from phonetic perception to syntactic and semantic integration.
    Compared with EEG, MEG generally provides better spatial localization because magnetic fields are minimally distorted by skull and scalp tissues, reducing uncertainties associated with EEG inverse problem [92]. This combination of high temporal resolution and improved (though still limited) spatial accuracy makes MEG particularly suitable for studies requiring precise spatiotemporal characterization of language processes, such as phonetic-to-lexical conversion and syntactic processing [208, 210]. However, MEG is primarily sensitive to tangential cortical sources (e.g., in sulci such as the superior temporal and intraparietal sulci) but is largely insensitive to radial sources (e.g., gyral crowns of the inferior frontal gyrus, angular gyrus, and motor cortex) and to subcortical structures. Source localization also remains an ill-posed inverse problem, and spatial resolution, though better than EEG, remains limited to the order of millimeters to centimeters.
  4. iv.
    ERP (Event-Related Potential): ERP is not an independent measurement technique but an analysis method applied to EEG recordings [149]. EEG signals are time-locked to stimulus onset and averaged across repeated presentations, extracting stimulus-related neural responses while reducing ongoing background activity and noise [128]. ERP therefore inherits EEG’s millisecond temporal resolution and provides precise temporal information about the time course of language processing, from early sensory analysis to lexical, semantic, and syntactic integration.
    Many ERP components have been identified in language research, among which N400 and P600 are the most widely studied. The N400 is a negative deflection peaking approximately 400 ms after stimulus onset and is strongly associated with semantic processing; its amplitude increases when words are semantically incongruent or unpredictable in context. The P600 is a positive deflection peaking around 600 ms and is primarily linked to syntactic processing; its amplitude increases in response to syntactic violations or elevated structural complexity [229, 178, 127].
  5. v.
    fNIRS (functional Near-Infrared Spectroscopy): fNIRS is a non-invasive optical imaging technique that measures brain activity through changes in the absorption of near-infrared light by oxygenated and deoxygenated hemoglobin [197, 235]. Near-infrared light emitted from scalp-mounted optodes penetrates superficial cortical tissue and is differentially absorbed by oxyhemoglobin and deoxyhemoglobin. Detectors capture diffusely reflected light, and concentration changes are estimated using the modified Beer–Lambert law, allowing indirect inference of neural activity. Like fMRI, fNIRS relies on hemodynamic responses associated with underlying neural activation.
    fNIRS offers several practical advantages. The equipment is portable and does not require shielded environments, enabling experiments in more natural communication settings. It is also generally less sensitive to motion artifacts than fMRI and, in some contexts, more robust than EEG, making it particularly suitable for studies involving infants, children, or patients with language disorders, as well as for tasks involving overt speech or facial movements [242]. Its temporal resolution lies between fMRI and EEG/MEG, allowing tracking of sub-second hemodynamic changes during tasks such as lexical access and sentence comprehension. Spatial resolution is typically on the order of centimeters, sufficient for localizing broad cortical language regions, though not for resolving fine-grained functional organization [71, 211].
    Limitations include lower spatial resolution than fMRI and limited penetration depth, which precludes reliable measurement of subcortical structures. fNIRS signals are also influenced by individual physiological factors such as scalp thickness, hair properties, and superficial (extracerebral) blood flow, requiring careful experimental design to ensure comparability across participants [137, 191]. Although whole-head coverage is now technically feasible with modular high-density arrays, it remains less common than localized cortical applications due to increased demans on hardware, setup time, and data processing, fNIRS is therefore more suitable for targeted investigations of cortical regions rather than comprehensive brain mapping [211].
  6. vi.
    ECoG (Electrocorticography): ECoG is an invasive neurophysiological technique that records electrical potentials directly from the exposed cortical surface using subdural electrode grids or strips placed during surgery [100]. Unlike scalp EEG, ECoG avoids signal attenuation and distortion by the skull and scalp, yielding high signal amplitude, excellent spatial resolution, and sub-millisecond temporal precision [112]. These characteristics make it uniquely capable of resolving fine-grained, region-specific cortical dynamics underlying language processing, including articulatory planning, phonemic encoding, lexical access, and syntactic binding. ECoG is widely used in presurgical mapping for patients with epilepsy or brain tumors, allowing direct localization of essential language areas such as Broca’s region, superior temporal gyrus, and inferior parietal cortex with high clinical and scientific reliability [53]. High-frequency broadband power modulations in ECoG are particularly sensitive to local neuronal population activity and correlate strongly with task-specific language computations. Key limitations include its invasive nature, restricted clinical eligibility, limited spatial coverage determined by electrode placement, and inherent ethical constraints that preclude use in healthy participants.
  7. vii.
    SEEG (Stereo‐Electroencephalography): SEEG is an invasive, stereotactic technique that uses multiple thin depth electrodes implanted through small burr holes to record electrical activity from deep brain structures and cortical regions, including those buried in sulci, while traversing white matter tracts. This enables precise three-dimensional mapping of the epileptogenic zone [263]. Unlike ECoG, which samples from the cortical surface, SEEG provides three-dimensional access to both cortical and subcortical targets critical for language control, semantic memory, and pragmatic integration [6]. It combines sub-millisecond temporal resolution with precise three‐dimensional localization, enabling characterization of spatiotemporal dynamics with language networks during both overt and covert processing [269]. SEEG is extensively used in epilepsy presurgical evaluation to map language-related networks and identify epileptogenic zones while preserving essential functions. Similar to ECoG, SEEG is restricted to clinical populations, involves surgical risk, and offers limited whole-brain sampling density [183]. Nevertheless, its unique ability to record from deep and hidden structures makes SEEG irreplaceable for investigating subcortical‐cortical circuits supporting language comprehension, production, and monitoring.
3.3 Word Embeddings
Numerical representations of language provide a critical bridge between linguistics, computer science, and neuroscience. Word embeddings, also referred to as distributed word representations, have become a foundational tool in Natural Language Processing (NLP) and in computational studies of neural language mechanisms [158]. The central idea is to map discrete words onto a continuous, high-dimensional vector space, where semantic similarity is reflected in geometric relations among vectors. This mapping enables symbolic linguistic units to be processed quantitatively by computational models. Unlike traditional one-hot encoding, which merely indicates word identity and captures no semantic relationships, word embeddings represent words through distributed, low-dimensional numerical features learned from distributional patterns in text. These representations capture rich semantic and syntactic associations and have significantly improved performance across a wide range of language tasks, from lexical similarity judgments to machine translation and brain encoding models [93].
Word embeddings are typically learned using machine learning or deep learning models and are grounded in the distributional hypothesis, which holds that word meaning is derived from patterns of contextual usage [207]. Based on training objectives and model architectures, mainstream approaches fall into two broad categories. The first comprises static embedding models such as Word2Vec and GloVe. Word2Vec learns word vectors by predicting local context using either the Skip-gram or CBOW architecture, while GloVe integrates global word-co-occurrence statistics with local context information. Both methods generate fixed vector representations that capture semantic similarity and analogical relations among words [186, 40]. The second category consists of context-dependent embeddings produced by modern Transformer-based pre-trained language models. These models dynamically produce context-sensitive representations, addressing polysemy (a limitation static embeddings cannot resolve) and substantially enhancing semantic representation quality [120, 1]. All such approaches rely on large-scale unlabeled corpora and learn patterns of word co-occurrence that place semantically related words in close proximity within the vector space, providing effective foundations for a wide range of downstream language tasks.
Researchers have extended lexical embeddings to larger linguistic units such as phrases, sentences, and discourses to represent complex structures numerically. Existing approaches can be grouped into three types. The first uses simple arithmetic composition, typically averaging or summing constituent word vectors [161]. These methods are computationally efficient and were widely used in early sentence-level tasks. However, by treating words as independent units, they ignore word order, semantic roles, and syntactic structure. As a result, expressions with opposite meanings but identical vocabularies may receive similar representations, limiting semantic precision [180, 41]. The second category employs structured composition methods, such as tensor product representations [222]. These approaches combine word embeddings through tensor operations to encode syntactic and combinatorial relations explicitly, preserving hierarchical and relational information within linguistic structures. Although this framework offers principled structured representations, high-dimensional tensor operations significantly increase computational cost and limit large-scale applications [160, 159, 42]. The third category includes task-driven dynamic representation methods that learn sentence encoders through supervised downstream tasks [46]. The assumption is that optimizing task performance encourages models to encode semantic and structural information automatically. However, empirical studies show that models often exploit superficial statistical patterns rather than genuine semantic structure, leading to reduced robustness and poor generalization under syntactic variation or domain shifts [167, 12].
Word embeddings play a central role for three main reasons. First, they connect computational linguistics with neurolinguistics [132]. The distributional principle underlying embeddings is broadly consistent with the brain’s distributed representation of lexical semantics, and similarities in embedding space often correspond to similarities in neural activity patterns measured with techniques such as fMRI and MEG [162]. Understanding these representations therefore facilitates direct comparison between computational and neural semantic structures. Second, embeddings provide the numerical input foundation for modern language models. Natural language is symbolic and discrete, and must be transformed into continuous numerical representations before it can be processed by neural networks. From early static embeddings to the context-dependent representations of LLMs, embedding quality strongly influences model performance [58]. Since later sections examine deep learning models for neural simulation, reviewing embedding principles clarifies how linguistic information enters computational models. Third, research on extending embeddings to complex expressions informs studies of neural representations of higher-level linguistic structure [8]. This work addresses not only lexical processing but also sentence- and discourse-level representations. The strengths and weaknesses of existing extension methods, particularly regarding syntactic preservation and semantic robustness, mirror challenges in modeling hierarchical neural language processing. Reviewing these methods therefore provides conceptual guidance for investigating how the brain encodes complex linguistic structures.
In summary, word embeddings and their extensions provide the core framework for numerical language representation. They have driven rapid progress in NLP while also offering essential tools and theoretical insights for interdisciplinary research on the neural mechanisms of language.
3.4 The Evolving Language Models
Language models are central tools at the intersection of computational neuroscience and linguistics. By learning statistical regularities from large text corpora, they extract latent semantic and syntactic information and generate outputs that approximate human language conventions. Their core computational objective is to predict lexical sequences based on preceding or surrounding context.
Research on language models is important on both theoretical and practical grounds. Theoretically, their development parallels efforts to understand the neurocognitive mechanisms underlying human language processing. Evaluating how well these models capture contextual dependencies and hierarchical structure provides quantitative benchmarks for studying the computational operations of the brain’s language network. Practically, language models form the backbone of many NLP applications, including machine translation, text generation, semantic analysis. Improving their alignment with human cognition is therefore widely regarded as an important step toward more general and robust artificial intelligence systems.
Table 3: Characteristic comparison of the evolving language models.
Model Type Core Architecture / Mechanism Linguistic Hierarchy Modeling Capability Representative Models Limitations
n-gram Statistical co-occurrence probability No hierarchical structure Trigram model Poor generalization ability
PCFG Explicit modeling of syntactic rules Supports long-distance dependency modeling Probabilistic PCFG Weak rule generalization
RNNs Sequential recurrence, gating mechanisms Implicitly captures long-distance dependencies LSTM, GRU Low serial computing efficiency
Transformer/LLMs Attention mechanism, parallel computation Supports ultra-long context processing GPT, BERT Lack of biological structural constraints
LLM-based Agents LLMs coupled with planning, tools, and feedback loops Multi-level linguistic reasoning via interaction AutoGPT, LangGraph Limited interpretability and long-horizon stability
The development of language models has progressed from statistical approaches to neural architectures, with many design principles influenced by insights from neuroscience. Major model families include n-gram models, Probabilistic Context-Free Grammars (PCFG), Recurrent Neural Networks (RNNs), Transformer architectures, pre-trained LLMs and LLM-based Agents. Successive generations have improved contextual modeling and representation capacity, ultimately enabling the emergence of LLMs. Their main characteristics are summarized in Table 3, and briefly introduced below.
  1. i.
    n-gram: n-gram models represent the earliest form of language modeling. They predict a word based on the co-occurrence statistics of a fixed number of preceding words. For example, a trigram model conditions predictions on the previous two tokens [232]. Although computationally simple and efficient, these models cannot represent hierarchical structure or capture long-distance dependencies, which limits their ability to model complex linguistic phenomena.
  2. ii.
    PCFG: Probabilistic Context-Free Grammars explicitly model hierarchical syntactic structure and can, in principle, capture certain long-distance dependencies. However, both PCFG and n-gram approaches suffer from limited generalization: they cannot exploit semantic similarity across words or structural similarity across contexts. As a result, they have been largely replaced by neural network models, which learn distributed representations and achieve substantially better generalization performance [110].
  3. iii.
    RNNs: Recurrent Neural Networks represent the first generation of neural language models. They encode contextual information by mapping words into embedding spaces and updating hidden states sequentially as each new input is processed [67]. Although RNNs do not explicitly encode hierarchical structure, their recurrent connections allow them, in principle, to capture contextual dependencies, including long-distance relations. In practice, however, RNNs suffer from vanishing gradients, which limits their ability to learn very long-range dependencies. Long Short-Term Memory (LSTM) networks are a widely used variant of RNNs designed to mitigate the vanishing gradient problem through gating mechanisms. This architecture substantially improves long-range dependency modeling and became the dominant approach in early neural language modeling systems [102].
  4. iv.
    Transformer Model: Transformer architectures have become the dominant framework for language modeling. Unlike RNNs, which process inputs sequentially, Transformers employ attention mechanisms that allow each token to interact with all others simultaneously [239]. This enables parallel computation and efficient modeling of global context. While Transformers do not explicitly encode linguistic hierarchy, dynamic attention weighting and scalable parallel computation allow them to model long-range dependencies effectively in practice. These properties have also enabled large-scale training on massive corpora, forming the technical foundation for modern LLMs.
  5. v.
    Pre-trained LLMs: Pre-trained LLMs represent the current framework in NLP [26]. Their core training strategy follows a two-stage pipeline: large-scale self-supervised pre-training on massive unlabeled corpora, followed by fine-tuning or instruction adaptation for downstream tasks. Pre-training allows models to acquire syntactic patterns, semantic associations, and long-range dependencies from large textual datasets, while fine-tuning adapts these capabilities to specific applications. Owing to large parameter scales and extensive training data, these models significantly surpass earlier approaches in language understanding and generation. A key advantage is the separation between general language learning and task adaptation, enabling transfer across tasks without retraining from scratch.
    Current LLM systems can be broadly categorized based on architecture and training objectives:
    • Autoregressive LLMs: These models generate text sequentially from left to right by predicting each token based on preceding context. The GPT series is the most prominent example, with parameter scales growing from millions to trillions and training corpora spanning diverse web and literary sources [199, 1, 26, 200, 179]. Such models excel at text generation, dialogue, and long-form continuation tasks. Related systems include ERNIE Bot and the GLM family. Their primary strength lies in fluent generation, although strictly unidirectional modeling limits access to full bidirectional context [228, 63, 267, 103].
    • Autoencoding Pre-trained Models: These models use masked token prediction and thus implement bidirectional language modeling. BERT is the most influential example, achieving major improvements in tasks requiring semantic understanding, including classification and reading comprehension [58]. Variants such as RoBERTa, ALBERT, and SpanBERT further refine training strategies and architectures. However, such models are less suitable for long-form text generation [144, 131, 111].
    • Encoder–Decoder Hybrid LLMs: These models combine bidirectional encoding with autoregressive decoding, balancing semantic understanding and generation. Representative systems include T5 and multilingual variants such as mT5. They perform well on tasks requiring both comprehension and generation, such as translation and summarization [200, 201, 260].
    • Multimodal Ultra-large LLMs: Recent systems extend text-only models to incorporate visual, audio, and video inputs, enabling multimodal reasoning and generation. Models such as Gemini integrate cross-modal information and expand capabilities from text processing to multimodal cognition and reasoning, marking an important direction toward more general AI systems [230, 231, 45].
  6. vi.
    LLM-based Agents: LLM-based agents represent a higher-level application paradigm that extends the capabilities of pre-trained large language models beyond passive language understanding and generation [150]. Their core design integrates autonomous task planning, multi-tool coordination, long-horizon interaction, and closed-loop self-reflection, enabling models to operate in dynamic environments and pursue goal-directed behaviors. By coupling general-purpose language models with execution logic and environmental feedback, LLM agents partially overcome the limitations of standalone LLMs in complex, open-ended tasks that require sustained reasoning and iterative decision-making. In linguistics, this paradigm provides a novel computational framework for addressing problems characterized by high theoretical abstraction and empirical complexity, thereby facilitating closer integration between data-driven modeling and theory-driven analysis [227].
    From an application perspective, several representative research directions have begun to emerge:
    • Applied Linguistics and SLA: LLM agents can support human-in-the-loop systems for language assessment and second language acquisition (SLA) tutoring [91]. They can generate proficiency tests, evaluate open-ended learner responses, and provide personalized feedback, while longitudinal interaction yields large-scale empirical data relevant to interlanguage development [215, 86].
    • Corpus Linguistics: To reduce costly and inconsistent manual annotation, LLM agents can orchestrate syntactic parsers, semantic resources, and clustering algorithms for multi-level corpus annotation, extending beyond part-of-speech and syntax to semantic roles, pragmatic functions, and discourse structure [163, 264].
    • Historical and Cognitive Linguistics: Multi-agent interaction systems offer controllable simulations of language origins, grammatical change, and the emergence of communicative conventions. By manipulating interaction environments and communicative pressures, researchers can test hypotheses about least effort and communication efficiency [192, 165].
    • Pragmatics: Leveraging advanced contextual modeling and reflective reasoning, LLM agents can be used to probe higher-order pragmatic phenomena (e.g., irony, metaphor, euphemism, presupposition) under controlled experimental settings, and to compare model-based pragmatic inference with human data across cultural and social contexts [108, 209].
    • Endangered Language Documentation: In low-resource settings, LLM agents show promise for assisting phonological rule induction, syntactic paradigm extraction, and transcription or normalization of spoken corpora, thereby supporting the construction of digital archives and long-term preservation efforts [224, 268].
    Despite these promising developments, applications of LLM agents in linguistic research remain at an exploratory stage, and no unified research paradigm or standardized technical framework has yet emerged. Current challenges can be summarized along three main dimensions. First, the interpretability of linguistic reasoning processes remains limited, as the internal mechanisms underlying grammatical induction and pragmatic judgment are largely obscure. Second, cultural and contextual generalization is insufficient, particularly for language phenomena with strong regional or historical specificity, where training data biases can be pronounced. Third, long-horizon interaction often suffers from consistency degradation, making it difficult to maintain stable reasoning strategies and research objectives over extended tasks. Future research should therefore prioritize the development of linguistically informed agent architectures that integrate domain-specific linguistic tools, enhance pragmatic reasoning and cultural adaptation, and incorporate explainable artificial intelligence techniques. Such efforts are essential for narrowing the gap between data-driven modeling approaches and theory-driven linguistic research, and for advancing linguistics toward an interdisciplinary framework that combines quantitative simulation with qualitative validation.
3.5 Neural Coding
Neural coding is a central analytical framework for linking computational models with neural mechanisms of language processing. Together with neural decoding and representational similarity analysis, it provides standardized tools for quantifying correspondence between computational representations and neural activity patterns [124, 168]. Broadly defined, neural coding methods construct mappings between stimulus representations and neural responses using statistical or machine learning models, allowing neural activity elicited by stimuli to be predicted computationally [162]. In language research, stimuli include not only spoken or written inputs but also internally generated semantic and syntactic representations. These stimuli are typically represented numerically using language models trained on large corpora, enabling quantitative comparison with neural recordings [115].
Neural coding analyses generally follow three steps. First, stimulus representations are paired with recorded neural signals to train mapping models, often using regression or shallow neural networks. Second, trained models predict neural responses to previously unseen stimuli. Third, predicted and measured responses are compared using metrics such as correlation or prediction error to evaluate model performance. Higher predictive accuracy indicates stronger alignment between model representations and neural mechanisms of language processing [106, 169].
Neural coding plays a critical role in evaluating computational language models. Standard NLP metrics such as perplexity or task accuracy do not directly reflect biological plausibility. Neural coding instead measures whether model representations correspond to neural activity patterns, providing evidence about whether models capture cognitively relevant features rather than merely exploiting surface statistics. It therefore offers a quantitative bridge between computational modeling and neural mechanisms of language processing [162, 69].
Neural coding is also essential for assessing alignment between LLMs and human neural language systems [81]. Although LLMs achieve impressive performance, it remains unclear whether their internal representations match brain mechanisms. Neural coding provides a quantitative test: if model-derived representations accurately predict neural responses, representational alignment is supported. Conversely, weak prediction suggests that models rely on processing strategies distinct from human cognition. Neural coding thus functions both as an evaluation tool for cognitive plausibility and as a guide for developing more brain-aligned language models [258].
Figure 3: Computational Neuroscience Workflow of Each Component Collaboration in the Linguistic Case Study.
3.6 “Model–Brain Alignment” Workflow from Computational Neuroscience in Linguistic Case
The methods introduced above operate as an integrated research framework rather than independent components (shown in Fig. 3). To illustrate their interaction, we consider a representative research scenario in which one tests whether LLM representations can predict neural activity during sentence reading. In this representative scenario, the framework can be unpacked as follows:
  1. i.
    Measurement Methods $\rightarrow$ Neural Data Acquisition
    $\circ$ Function: Techniques such as fMRI, EEG, MEG, ERP, fNIRS, ECoG and SEEG, record neural responses during language processing, providing complementary temporal, spatial, and representational information.
    $\circ$ Role in alignment: These recordings serve as empirical benchmarks against which model predictions are evaluated.
  2. ii.
    Word Embedding $\rightarrow$ Linguistic Representation Basis
    $\circ$ Function: Linguistic units including words, phrases, and sentences are mapped to continuous vector representations, enabling computational processing of semantic information.
    $\circ$ Role in alignment: Embedding quality influences whether model representations capture semantic structures comparable to neural representations. Similarity between embedding spaces and neural activity patterns is often examined as preliminary evidence of alignment.
  3. iii.
    Neural Coding $\rightarrow$ Model–Brain Mapping
    $\circ$ Function: Neural coding constructs predictive mappings from model representations to neural signals, allowing quantitative evaluation of correspondence between internal model states and brain activity.
    $\circ$ Role in alignment: It directly evaluates whether model representations predict neural responses, providing an operational criterion for functional alignment.
  4. iv.
    Language Models $\rightarrow$ Computational Reference Frame
    $\circ$ Function: Language models, ranging from statistical models to Transformer-based systems, learn internal representations of linguistic structure from large corpora.
    $\circ$ Role in alignment: These representations provide candidate computational spaces whose correspondence with neural representations can be tested using neural coding or representational similarity analysis.
To clarify how these components interact, we outline a hypothetical study examining whether LLM representations predict neural activity during sentence comprehension. Assuming the objective is to test whether LLM-derived representations predict neural responses in specific brain regions while participants read sentences. The procedure can be summarized as follows:
  1. i.
    Measurement Methods: A participant reads sentences presented sequentially while undergoing fMRI scanning. Whole-brain BOLD responses are recorded, yielding time-resolved spatial activity maps. These measurements serve as the empirical neural responses used for alignment evaluation.
  2. ii.
    Language Models & Word Embeddings: The same sentences are processed by a pre-trained LLM. Words are first converted into embeddings, transforming discrete tokens into continuous representations. These embeddings pass through multiple Transformer layers, where increasingly abstract linguistic features are extracted. Lower layers typically encode local lexical and syntactic information, whereas deeper layers tend to reflect more global semantic structure. For each sentence, activation vectors at different layers—often taken at the final token position—are extracted as candidate computational representations. Each layer thus provides a hypothesis about how sentence information may be encoded computationally.
  3. iii.
    Neural Coding: Sentence data are partitioned into training and test sets. For a given model layer, activation vectors from training sentences serve as inputs, while neural responses from a selected brain regions of interest (ROI) serve as outputs. A regression model is trained to map model representations to neural activity. The trained mapping is then applied to activation vectors derived from test sentences to predict neural responses. Prediction quality is evaluated by comparing predicted and measured neural signals, commonly using correlation metrics such as Pearson’s correlation $r$. High predictive accuracy indicates that representations from that model layer align with neural responses in the ROI. Repeating this analysis across layers reveals which representational levels correspond most closely to specific brain regions. For example, stronger correspondence between middle model layers and semantic brain regions would support the hypothesis that intermediate model representations capture semantic structure. Neural coding thus provides a quantitative mechanism for evaluating alignment, replacing qualitative claims of similarity with measurable predictive performance.
This example illustrates the complementary roles of each component: measurement methods provide empirical neural responses; language models and embeddings supply candidate computational representations; neural coding quantitatively evaluates which representations best correspond to neural activity. Demonstrating alignment not only tests the cognitive plausibility of language models but also allows computational models to serve as analytical tools for interpreting neural processing mechanisms, enabling hypotheses about how linguistic information may be encoded in the brain.
4 Foundational Insights and Framework Shifts in Language-Brain Computational Research
Existing research commonly introduces or adopts language processing models to examine the relationship between language and the brain from multiple perspectives. One major point of differentiation concerns model design motivation. Some studies employ architectures inspired by cognitive or linguistic theories, explicitly aiming to simulate particular aspects of human language processing. Other studies rely on engineering-oriented LLMs originally developed for NLP applications, which nonetheless provide high-dimensional representations and performance metrics useful for analyzing linguistic processes. Although LLMs largely originate from engineering practice, their conceptual foundations relate to parallel distributed processing (PDP), a framework initially proposed to explain neural information processing in the brain [205]. With rapid recent development, LLM-based analyses have increasingly contributed to understanding how language processing in artificial systems relates to neural mechanisms in humans.
4.1 Early Model Exploration of Predictability and Structural Features
Early investigations into the relationship between language processing and brain activity focused primarily on predictability, sequential and structural information, and neural encoding and decoding frameworks. These studies largely conceptualized language comprehension as an online, temporally unfolding process, in which moment-by-moment expectations and updates could be quantitatively linked to neural signals recorded with noninvasive techniques such as EEG, MEG, or fMRI [23]. At the same time, invasive intracranial electrophysiological recordings, particularly ECoG, complemented these approaches by providing high temporal and spatial resolution access to cortical dynamics during naturalistic speech perception and production [173].
A representative line of work employs $n$-gram language models and related variants. These models estimate word surprisal from conditional probabilities in context, which are then used to predict neural responses [88, 253]. Owing to their simplicity and broad applicability, such methods are well suited for naturalistic narrative studies. Empirical results show that surprisal reliably predicts neural responses associated with language processing, including modulation of the N400 component, supporting the view that predictive mechanisms operate continuously during comprehension [73]. However, these models primarily capture surface-level sequential regularities and make limited commitments to the hierarchical structure of language. Consistent effects of surprisal and predictability have also been observed during continuous speech processing, including in ECoG recordings where high-gamma activity in superior temporal and frontal regions tracks word-level expectations, reinforcing the role of predictive coding at fine temporal scales [164].
This limitation prompted subsequent work to move beyond purely sequential models by incorporating hierarchical syntactic structure. Researchers introduced phrase structure grammars, recursive syntactic models, or explicit syntactic representations to examine whether hierarchical dependencies improve explanation of neural data. Results indicate that hierarchical information contributes uniquely to explaining EEG and fMRI responses, suggesting that human language processing integrates both local lexical statistics and abstract structural dependencies [214, 21]. Related work further suggests that such hierarchical and combinatorial structure is reflected in distributed cortical responses during both speech perception and production, a pattern that is also evident in intracranial recordings with distinct temporal dynamics across temporal and frontal language areas [173].
Methodologically, the growing diversity of linguistic models—ranging from sequential surprisal-based approaches to hierarchical syntactic representations—necessitated a common analytical framework for systematic comparison. In this context, the neural encoding framework emerged as a central analytical tool. This approach uses internal representations from computational models to predict neural responses, or learns mappings from neural activity to linguistic features. It enables comparison across models and examination of how model architecture, training scale, or input representations affect prediction performance [251, 234]. Findings from encoding studies suggest that lexical semantics, syntactic structure, and semantic updating processes all contribute to model–brain correspondence. Such encoding approaches have proven particularly effective for speech data, as model-derived features can predict high-frequency cortical activity during naturalistic listening and speaking [164].
Complementary to encoding approaches, decoding frameworks reverse the mapping by treating neural activity patterns themselves as representations of linguistic content and training classifiers to distinguish stimuli or semantic categories directly from brain data [221]. This perspective highlights the discriminability and generalization properties of neural representations, though challenges remain regarding interpretability and construct validity. Decoding studies further demonstrate that phonetic, lexical, and semantic information can be reliably recovered from cortical activity during speech, with particularly clear evidence from ECoG recordings [184].
At the neural component level, questions about prediction, structure, and semantic updating converge most prominently on the N400, which has become a major focus connecting language models and neural signals. Studies modeling N400 responses using surprisal, semantic similarity, and semantic updating metrics demonstrate that multiple predictive factors can exert independent or complementary effects on amplitude [198, 148]. These findings suggest that the N400 reflects combined predictive and integrative processing across linguistic levels. Although the N400 is classically characterized in scalp EEG, converging intracranial evidence, including ECoG, links analogous semantic and predictive effects to localized cortical dynamics during speech comprehension [24].
In summary, early computational studies of language–brain alignment indicate that neural language processing is simultaneously sensitive to lexical predictability, hierarchical syntactic structure, and semantic updating. Sequential models, structural representations, and encoding and decoding frameworks capture complementary aspects of model–brain correspondence and together establish the theoretical questions and methodological foundation for later work incorporating deep learning–based models into neural alignment studies. Across recording modalities, these findings collectively underscore the continuity between computational models of language and the neural mechanisms supporting both written and spoken language processing.
4.2 LLM-Driven Framework for Neural Alignment
Compared with the gradual advances achieved by statistical language models and early neural architectures, the emergence of LLMs has substantially accelerated interdisciplinary research at the intersection of linguistics, neuroscience, and artificial intelligence, particularly in studies examining alignment between computational models and neural mechanisms of language processing [258, 3, 196]. Two developments largely account for this shift. First, large-scale self-supervised pre-training on massive corpora enables models to capture broad patterns of syntax, semantics, and language use. Second, deep neural architectures provide distributed representations capable of integrating long-range contextual information [151, 259, 244, 213]. Consequently, LLMs display behavioral performance resembling human processing across tasks involving contextual integration, syntactic analysis, semantic reasoning, and text generation. In several domains, including ambiguity resolution and anaphora interpretation, performance approaches or surpasses human benchmarks, making these models valuable tools for investigating computational aspects of language cognition [11, 249, 105].
More importantly, LLMs have increasingly become analytical instruments for studying neural language mechanisms [219, 262, 135]. Earlier work frequently relied on correlations between behavioral results and brain imaging data without an intermediate computational framework. LLMs now provide explicit representational systems linking linguistic input, behavioral outcomes, and neural responses. Recent studies systematically compare hierarchical representations of LLMs with human brain imaging data to infer organizational principles of language networks, temporal dynamics of processing, and neural coding strategies underlying semantic representation [146]. This trend has strengthened integration between linguistics and neuroscience, enabling analyses that jointly consider computational representations, behavioral responses, and neural activity. Accordingly, the remainder of this section surveys recent progress in LLM-driven alignment research across five dimensions: cross-modal representation, inter-brain neural coupling, hierarchical language computation, learning strategies and data distribution sensitivity, and language plasticity with shared representational spaces.
4.2.1 Language Modulation of Perception: Cross-Modal Representations (Vision & Music)
Whether language actively shapes human perception of the external world remains a long-standing and contested issue in cognitive science and linguistics, often discussed under the framework of linguistic relativity. Early cross-linguistic studies, including work on color categorization and laboratory label-learning experiments, suggested that linguistic labels may influence categorical judgments and perceptual sensitivity to attributes such as color, shape, and texture. However, consensus remains lacking regarding the strength and interpretation of these effects [252, 14, 250, 114]. Debates primarily concern three issues: the robustness and replicability of behavioral effects across populations and paradigms; the typically small effect sizes observed in many experiments; and the processing stage at which language exerts influence, whether during early perceptual encoding or only in later decision or memory stages. These unresolved issues have long complicated interpretation of language–perception interactions [79, 27, 233].
Recent advances in vision–language models, particularly Contrastive Language–Image Pre-training (CLIP), have introduced new computational tools for addressing these debates and pushed research into a new interdisciplinary phase [13, 33]. Through joint training on large-scale image–text pairs, such models learn unified representational spaces spanning visual and linguistic modalities. Their controllable training conditions allow partial disentanglement of linguistic and visual influences. Neuroimaging studies consistently report that representations derived from vision–language models exhibit higher correspondence with neural representations in high-level visual regions such as the Ventral OccipitoTemporal Cortex (VOTC) than purely visual models, and better predict responses to novel visual stimuli [261, 143].
Two principal interpretations of this similarity remain under discussion. One proposes that language exerts a direct modulatory effect on visual representations, reshaping VOTC representational geometry through neural pathways. The alternative common-cause hypothesis argues that language and perception share underlying relational structures, and model similarities simply reflect learning of general multimodal associations rather than language-specific effects. The central disagreement concerns whether language plays a causal role in shaping visual representations [261, 32]. To address this issue, several studies have used patients with disrupted language–vision connectivity as natural experiments. By examining how VOTC representations change following disconnection of language pathways, these studies provide evidence relevant to causal interpretations. Using diffusion tensor imaging to identify damage in white matter pathways linking language networks and VOTC, researchers compare neural responses during visual classification tasks between patients and healthy controls. Results indicate that pathway integrity significantly influences VOTC representational structure: greater disconnection is associated with representations more closely resembling purely visual models, while healthy individuals show representations incorporating linguistic relational structure [31]. Studies of sign language users further indicate that both spoken and signed languages can influence visual processing, suggesting modality-independent language effects. Although such AI–brain–lesion comparisons provide strong converging evidence, larger multi-center studies remain necessary for confirmation [68].
Cross-modal representational interactions also arise between language and other symbolic systems such as music [185]. While neural substrates underlying low-level musical processing are relatively well characterized, mechanisms supporting representation of high-level musical semantics, including genre and emotion, remain less understood [118]. Meta-analytic work suggests partial neural overlap between language and music processing, particularly in auditory–motor circuits involved in phonological and interval processing, while higher-level structural processing shows greater domain specificity [96]. The development of MusicLM and related text–music generation models offers new computational tools for studying musical semantic representation. These models learn mappings between textual descriptions and musical features, enabling quantitative comparison with neural responses. Recent studies using fMRI recordings during music listening demonstrate that intermediate and higher-layer MusicLM representations enable reconstruction of musical segments from neural signals and better predict auditory cortex responses than generic auditory models. Moreover, multimodal music–text embeddings show overlapping predictive regions within auditory cortices [57].
Current evidence supports two preliminary conclusions. First, cross-modal interactions between language and perceptual systems such as vision and music exist and depend on specific neural pathways. Second, multimodal large models partially capture these interaction regularities and serve as useful computational tools for studying semantic coding in perceptual systems. Nonetheless, several open questions remain. Individual differences in linguistic modulation of perception remain poorly understood; model–brain discrepancies persist, particularly in fine perceptual detail processing; and developmental mechanisms underlying shared representations between language and music are unclear. Future research combining larger and more diverse samples with refined model comparison frameworks will be necessary to clarify the mechanisms and boundary conditions governing language modulation of perceptual systems.
4.2.2 Inter-Brain Neural Coupling in Language Communication: Speaker-Listener Circuitry
The central function of language is to enable transmission of information and coordination of cognition across individuals. Understanding how speech production and comprehension dynamically interact during real communication has therefore become a central topic in language neuroscience [190]. Early research typically employed isolated experimental frameworks, such as monologue production or passive listening, to localize neural substrates of production and comprehension separately. These studies established that speech production primarily involves the sensorimotor cortex, supplementary motor area (SMA), and inferior parietal regions responsible for phonological planning and articulation, whereas speech comprehension is centered on the superior temporal sulcus (STS) and auditory cortex, which support speech perception and early parsing [98, 78]. However, single-participant and non-interactive designs cannot capture real-time interactions between speakers and listeners, leaving unresolved how distinct production and comprehension systems coordinate across individuals during conversation [225, 266, 203].
Recent advances in hyperscanning techniques and LLMs have provided complementary tools for addressing this problem. Hyperscanning allows simultaneous recording of neural activity from interacting individuals, enabling direct measurement of inter-brain correlations during communication. At the same time, LLMs enable quantitative modeling of conversational semantic content through contextual embeddings that capture both lexical meaning and long-range dependencies. This combination helps separate semantic contributions from low-level acoustic influences and supports identification of factors driving inter-brain neural coupling. As a result, joint use of hyperscanning and LLM-based representations has become an increasingly common framework for studying neural mechanisms underlying natural language interaction [81, 266, 2].
Studies using this framework have yielded several consistent observations. Anatomically, production and comprehension systems remain largely distinct: production-related activity concentrates in sensorimotor and parietal regions, whereas comprehension-related activity is centered in temporal auditory regions, with limited spatial overlap. Functionally, however, strong inter-brain coupling emerges during successful communication. Neural activity in a speaker’s production network predicts activity in a listener’s comprehension network, and this coupling is not explained by acoustic similarity alone. Instead, coupling strength increases when interlocutors share aligned semantic representations of conversational content, a pattern observed across languages and communication settings [266].
Further work shows that coupling patterns adapt to communicative demands in both spatial and temporal dimensions. Spatially, coupling extends beyond classical language areas to include regions associated with social cognition and mentalizing, such as the temporoparietal junction (TPJ), posterior cingulate cortex (PCC), and medial prefrontal cortex (mPFC). This indicates that communication involves not only linguistic decoding but also alignment of intentions and shared knowledge between participants. Temporally, coupling often exhibits predictive characteristics: although comprehension signals typically lag production signals overall, higher-level cortical regions sometimes show anticipatory alignment with the speaker’s intended meaning before key information is fully expressed. These findings provide empirical support for predictive processing accounts of language comprehension and parallel computational principles employed in modern LLMs [190, 175, 77].
Inter-brain coupling research carries both theoretical and methodological implications. Theoretically, it challenges views that treat production and comprehension as isolated processes, instead proposing communication as functional coupling between specialized systems across individuals [138]. Such mechanisms help explain communication success under noisy conditions and may illuminate communication impairments observed in disorders such as autism, aphasia, and schizophrenia, where altered coupling strength or instability has been reported [101, 64]. Methodologically, combining LLM-based representations with hyperscanning extends alignment research beyond single-brain analyses toward dynamic interaction, reinforcing the utility of computational language models in neuroscience.
Nevertheless, important questions remain. Individual differences in coupling dynamics are poorly characterized; cross-modal communication mechanisms remain underexplored; and the extent to which LLM representations capture pragmatic and interactional aspects of communication is still unclear. Future work combining multimodal hyperscanning, conversational LLMs, and more diverse participant populations will help refine spatiotemporal models of inter-brain communication and clarify boundaries of model–brain alignment.
4.2.3 Hierarchical Language Computation: Human Brain vs. LLMs
Language exhibits hierarchical structure, progressing from phonological and lexical processing to syntactic and semantic integration, thereby enabling efficient information transmission [36]. Neuroscientific research has consistently implicated core language regions such as the left inferior frontal gyrus (Broca’s area), superior temporal gyrus, and middle temporal gyrus in hierarchical processing. Broadly, Broca’s area contributes to syntactic construction and repair, whereas temporal regions support lexical-semantic activation and integration [76, 83]. Transformer-based LLMs similarly demonstrate strong performance on tasks involving hierarchical structures such as nested dependencies and long-distance agreement [141]. However, a central debate persists: whether LLMs genuinely implement brain-inspired hierarchical computation or merely reproduce similar behavioral outcomes through large-scale statistical learning.
To examine this issue, recent work has developed methods linking syntactic representations in models and neural activity patterns. The Hierarchical Frequency Probe approach, for example, uses frequency-domain analyses to identify populations encoding syntactic structure in both neural data and model representations [4, 60]. Comparative studies using multiple LLM families alongside fMRI recordings of subjects processing syntactically complex sentences reveal consistent findings: in models, syntactic information is predominantly encoded in intermediate and higher layers, whereas lower layers primarily capture lexical features. In human cortex, lower-level syntactic information is associated with anterior temporal regions, while higher-level syntactic processing engages Broca’s area and adjacent regions [76, 181]. Representational similarity analyses further indicate stronger correspondence between model representations and left-hemisphere language regions than with right-hemisphere homologues, supporting partial functional correspondence.
Temporal correspondence has also been investigated to address differences between biological sequential processing and model parallel computation. Using ECoG, researchers have recorded neural responses during natural story listening and compared time-resolved neural activity with representations from different model layers [81]. Results show systematic correspondence between processing stages: shallow model layers align with early neural responses, whereas deeper layers correspond to later neural activity, particularly within core language regions. Despite hardware-level differences between brains and models, hierarchical organization thus appears functionally mappable onto neural processing time courses [54].
Research has further extended to predictive processing over longer time scales. Human language comprehension relies heavily on prediction, and recent fMRI studies have compared neural responses during narrative listening with predictions generated by models of varying complexity [265, 245, 270]. Findings indicate hierarchical prediction mechanisms: superior temporal regions support local lexical predictions resembling shallow model layers or n-gram models, whereas default mode network regions contribute to global predictions over sentence and discourse scales, aligning more closely with higher-level LLM representations. Sparse update prediction models appear to better capture neural dynamics at discourse boundaries, suggesting multi-time-scale predictive organization. LLMs exhibit comparable hierarchical prediction patterns when processing long texts, further reinforcing parallels between model and brain computations.
Current evidence therefore suggests multi-dimensional correspondence between LLMs and neural mechanisms underlying hierarchical language processing. However, several open questions remain. Correspondence at pragmatic and discourse reasoning levels remains insufficiently explored; individual differences in hierarchical processing are poorly understood; and it remains unclear whether model hierarchies reflect genuine structural computation or statistical overfitting to text data. Future research combining multimodal imaging, diverse language tasks, and causal intervention methods will be necessary to clarify limits of model–brain correspondence and support both neuroscientific theory development and brain-inspired improvements in language models.
4.2.4 Learning Strategies & Data Distribution Sensitivity: Human Brain vs. LLMs
In neuroscience and artificial intelligence, a central question concerns how learning systems acquire rules and adapt to new tasks: do they rely primarily on weight-based learning (memorization) or contextual learning (inductive reasoning from limited examples)? The answer bears directly on our understanding of human learning and the foundations of intelligence in artificial systems [129, 74]. The emergence of strong in-context learning capabilities in LLMs has renewed interest in this debate. Without updating parameters, LLMs can rapidly adapt to new tasks by incorporating a small number of examples into prompts, behaviorally resembling rapid human inductive reasoning. However, direct empirical evidence remains limited regarding whether the underlying mechanisms are truly homologous, restricting deeper analysis of human-like intelligence and model optimization [126].
To address this issue, recent work introduced a standardized image–label associative learning framework to systematically compare humans and Transformer-based LLMs under controlled data distribution conditions [187]. Three distribution scenarios were designed: (1) highly diverse data, where image–label pairs rarely repeat and systems must infer general rules; (2) highly repetitive data, enabling performance gains through memory consolidation; and (3) mixed distributions that test adaptive strategy switching. Results showed strong convergence between humans and LLMs. Under diverse data, both relied primarily on contextual learning. Under repetitive data, both shifted toward memory-based strategies: humans consolidated memory traces, whereas LLMs encoded recurring patterns in internal representations, reducing computational cost. Under mixed conditions, both displayed flexible strategy use, dynamically selecting processing modes according to input structure. These findings provide behavioral and computational evidence that humans and LLMs share key statistical learning tendencies and support the human-like nature of in-context learning.
Important differences nevertheless emerged. Humans showed stronger resilience to distribution shifts, maintaining prior strategies while adapting to new ones and rapidly reinstating earlier strategies when distributions reverted. In contrast, LLMs exhibited strategy forgetting, requiring many new examples to recover previous processing modes. This divergence highlights differences in learning flexibility and in the interaction between memory stability and adaptive mechanisms.
Such comparative studies carry both theoretical and practical implications. Theoretically, they clarify that human learning involves coordinated interaction between long-term consolidation and contextual reasoning, enabling flexible adaptation while preserving stable knowledge structures [206, 59]. Practically, these findings suggest directions for improving in-context learning in LLMs, for example through strategy memory mechanisms or adaptive responses to distribution shifts, thereby bringing model behavior closer to human learning patterns [202, 166].
Several open questions remain. Existing work has largely focused on simple associative tasks, leaving distribution sensitivity in complex and naturalistic language learning—particularly for long-range dependencies, compositional structure, and pragmatic reasoning—poorly understood. Individual differences in human learners and performance variability across models and training regimes are also underexplored, limiting insight into the generalizability of learning strategies. Moreover, the neural mechanisms underlying these differences remain unclear due to the limited integration of brain imaging, cognitive modeling, and model interpretability approaches. Future research should therefore adopt more complex, naturalistic tasks; combine behavioral experiments with multimodal neural measurements; and recruit larger, more diverse participant samples. In parallel, systematic model analyses—including ablations, architectural comparisons, and training data manipulations—are needed to identify the computational sources of learning differences. Integrating these directions will help clarify divergences between human and LLM learning strategies and advance both language neuroscience and brain-inspired artificial intelligence.
4.2.5 Language Plasticity, Individual Differences & Shared Representations
In natural communication, the human language system must balance plasticity and stability, enabling adaptation across speakers and contexts while preserving consistent linguistic representations [125, 85]. On the one hand, listeners must rapidly adapt to accents, dialects, and contextual variations; on the other, phonological and semantic categories must remain stable to ensure reliable communication [123]. For example, listeners can quickly adapt to unfamiliar accents without permanently altering established phonological categories [43]. Although this adaptive stability is well documented behaviorally, its neural implementation remains incompletely understood, particularly regarding how plasticity and stability are balanced across processing stages [155, 17].
Recent EEG studies have addressed this issue using controlled manipulations of acoustic cues, such as fundamental frequency (F0) and voice onset time (VOT), to simulate unfamiliar accent conditions [145]. By degrading cue reliability, researchers examined how listeners adapt during word recognition while recording neural responses. Analyses of early perceptual components (N1, P2) and later semantic processing (N400) revealed a mechanism of selective cue downweighting. Rather than restructuring phonological categories, listeners reduced reliance on unreliable cues while maintaining processing of stable cues. Importantly, cue reweighting occurred at early perceptual stages without altering later semantic processing, thereby preserving representational stability. These findings clarify how the brain achieves adaptive flexibility without compromising core representations and offer potential guidance for improving accent adaptation mechanisms in LLM-based systems.
Parallel challenges arise from individual variability and limited sample sizes in language brain mapping, particularly in high-resolution techniques such as ECoG [27, 142]. Anatomical and connectivity differences across individuals complicate extraction of common neural patterns, while clinical constraints limit participant numbers. To mitigate these issues, recent work introduced the Shared Response Model (SRM) to align neural data across individuals within a shared low-dimensional representational space [15]. Using ECoG recordings during story listening and reading tasks, researchers applied SRM to standardize neural responses before comparing them with hierarchical representations in LLMs. Results showed that despite anatomical variability, aligned neural representations exhibited strong cross-individual consistency during language processing. These shared representations extended beyond classical left-hemisphere language regions and were also observed in right-hemisphere homologous areas, indicating bilateral contributions to language processing.
Further validation demonstrated practical benefits: neural encoding models trained in the shared space achieved improved prediction of unseen individuals’ brain activity, and semantic decoding performance substantially exceeded that of unaligned approaches. This work provides both a technical solution to data scarcity and individual variability and empirical evidence for shared neural coding principles underlying language processing. The resulting shared representational space also supports clinical translation, particularly in Brain–Computer Interface (BCI) development for language-impaired patients [223, 218, 254, 156].
Nevertheless, open questions remain. Mechanisms underlying individual differences in adaptive language plasticity remain poorly characterized, and the generalization limits of SRM in complex scenarios such as bilingual or dialectal processing require further testing. Moreover, correspondence between shared neural representations and hierarchical representations in LLMs remains insufficiently explored beyond semantic levels. Future studies should integrate multimodal imaging, larger and more diverse participant cohorts, and complex language tasks to refine models of language plasticity, improve cross-individual alignment techniques, and deepen brain–model correspondence analyses, thereby advancing both theoretical and clinical applications.
5 Discussion
5.1 Current Limitations
Despite substantial progress at the intersection of linguistics and neuroscience, recent advances reveal several persistent limitations that constrain current understanding and call for targeted future developments:
  1. i.
    Mechanistic interpretation remains limited by black-box mapping, and functional alignment does not imply structural homology. Most current studies evaluate correspondence between LLM representations and brain activity using methods such as Representational Similarity Analysis (RSA) and neural encoding models. However, correlation-based alignment demonstrates only functional similarity and does not establish shared computational mechanisms. Existing evidence suggests that similarities between LLMs and human language processing largely reflect functional fitting rather than mechanistic equivalence. Human language acquisition is grounded in multimodal interactive experience, whereas LLMs learn from unimodal textual data, limiting the explanatory power of alignment results regarding why model representations predict neural responses [48]. For example, although model attention mechanisms resemble human selective attention at a functional level, the former relies on parallel weight computation, whereas the latter depends on dynamically regulated neural circuits and neuromodulatory processes. No structural correspondence has yet been established, leaving open the possibility that observed alignment reflects statistical fitting rather than mechanistic simulation.
  2. ii.
    Technical constraints limit research depth and hinder real-time interactive verification. In studies combining LLMs with speech/language BCIs, current information transmission rates remain insufficient for efficient interaction between neural signals and model parameters, making direct model–brain mapping difficult. Neural recordings also suffer from low signal-to-noise ratios and strong individual variability, challenging accurate decoding of dynamic correspondences between neural activity and model representations. Cross-individual alignment methods such as SRM extract group-level regularities but do not preserve individual neural coding patterns [50]. As a result, most research remains limited to offline correlation analyses, preventing real-time closed-loop experiments and restricting investigation of dynamic language processing mechanisms.
  3. iii.
    Research coverage remains limited in both data diversity and task design. LLM training data are dominated by standardized written text, with limited representation of spoken language, dialectal variation, and multilingual mixing. Consequently, alignment studies may fail to capture neural mechanisms underlying natural communication. Experimental frameworks are also largely restricted to passive tasks such as reading or story listening, with comparatively little work examining active conversational interaction, pragmatic reasoning, or communicative intent transmission, despite these processes constituting the core function of language. In addition, neuroimaging studies predominantly involve healthy young adults, while data from older populations and individuals with language-related disorders such as aphasia or autism remain scarce. These biases limit generalizability and reduce potential clinical impact.
  4. iv.
    Evaluation frameworks remain narrow, and measures of brain alignment are incomplete. Current alignment studies typically rely on neural prediction accuracy as the primary metric, yet this single measure cannot fully characterize brain-inspired processing. Models may achieve high prediction accuracy through statistical shortcuts rather than genuine replication of hierarchical linguistic computations [48]. For instance, brain activity during syntactic processing may be predicted using surface lexical statistics rather than structural parsing, resulting in strong predictive performance but weak mechanistic correspondence. Furthermore, evaluation frameworks rarely incorporate human-specific cognitive properties such as language plasticity, long-term memory consolidation, or adaptive strategy switching, making it difficult to assess fundamental differences between LLMs and human language systems.
In summary, although interdisciplinary research has established productive connections between LLM development and language neural studies, substantial limitations remain in mechanistic interpretation, technical feasibility, ecological validity, and comprehensive evaluation. Future progress requires tighter integration across neuroscience, linguistics, and artificial intelligence, combined with technical innovation, richer experimental scenarios, and more comprehensive evaluation frameworks. Such advances are necessary both for clarifying model–brain correspondence and for promoting the development of brain-inspired language models alongside a deeper understanding of neural language mechanisms.
Figure 4: Correspondence between Current Limitations and Future Research Perspectives in the Study of Neural Mechanisms of Language.
5.2 Future Research Perspectives
Despite rapid progress in applying large language models to computational neuroscience, fundamental limitations remain in mechanistic interpretability, ecological validity, and translational applicability. Existing studies are still dominated by correlational analyses and static modeling paradigms, which restrict their ability to capture the dynamic, causal, and cognitively grounded nature of human language processing. Addressing these challenges requires future research to move toward integrated frameworks that jointly consider biological mechanisms, real-time interaction, data diversity, and evaluation standards.
In this context, four complementary research directions are particularly critical. First, strengthening structure–function matching aims to reduce the gap between artificial architectures and biological neural mechanisms through brain-inspired modeling and causal intervention. Second, constructing real-time closed-loop interaction systems emphasizes dynamic, bidirectional coupling between neural activity and language models. Third, expanding multimodal and multi-population linguistic neurodata seeks to improve ecological validity and generalizability. Finally, establishing multi-dimensional evaluation frameworks moves beyond neural predictability toward cognitive and mechanistic validation. Together, these directions define a coherent roadmap for advancing LLM-driven computational neuroscience (Fig. 4).
5.2.1 Strengthening Structure–Function Matching via Brain-inspired Modeling and Causal Interventions
Although Transformer-based LLMs can reproduce brain activity patterns at a functional or representational level, substantial differences remain in underlying computational mechanisms and energy efficiency. This structure–function mismatch limits deeper model–brain alignment and weakens claims of mechanistic correspondence. Strengthening structure–function matching therefore constitutes a foundational direction for future research.
Spiking Neural Networks (SNNs) provide a promising pathway toward improved biological plausibility. Inspired by event-driven communication in biological neurons, SNNs encode information through spatiotemporal spike patterns and naturally support temporally precise dynamics. Compared with conventional artificial neural networks, SNNs can achieve orders-of-magnitude improvements in energy efficiency while more faithfully reproducing neural signaling properties [243]. Recent studies have demonstrated the feasibility of integrating SNNs with large-scale language models through spiking attention mechanisms, hierarchical architectures, and model conversion techniques, enabling improved alignment without prohibitive performance loss [271, 248, 257]. The recently proposed “Shunxi 1.0” spiking large model further illustrates the potential of this approach for ultra-long sequence processing with high efficiency [182].
Beyond architectural similarity, future work should incorporate causal intervention experiments to validate structure–function correspondence. Targeted perturbations of model components, circuit-level ablations, and causal mediation analyses can be combined with neural stimulation or lesion data to test whether homologous structures in models and brains support comparable functional roles. Such causal validation moves beyond correlational alignment and enables stronger inferences about shared computational principles.
5.2.2 Constructing Real-time Closed-loop Language–Brain Interaction Systems
Most existing language–brain alignment studies rely on offline analyses of pre-recorded neural data, limiting their relevance to real-time cognition and interaction. Constructing closed-loop systems that dynamically couple neural activity and language models represents a critical future direction.
Low-power, event-driven SNN-based models are particularly well suited for real-time applications due to their computational efficiency and temporal precision. When integrated with online neural recording modalities, such models can support adaptive decoding and encoding of linguistic information, enabling continuous bidirectional interaction between brains and artificial systems. This capability is essential for capturing the dynamic nature of language processing and learning.
A major challenge in closed-loop systems is inter-individual variability in neural representations. Cross-individual alignment methods, such as Shared Response Models (SRM), can be incorporated to establish shared representational spaces across subjects, thereby reducing data requirements for new users. Combined with individual-specific fine-tuning, such hybrid general–specific frameworks may enable scalable and personalized real-time language–brain interaction systems, particularly for brain–computer interface applications.
5.2.3 Expanding Multimodal and Multi-population Linguistic Neurodata
Progress in computational language neuroscience is fundamentally constrained by the scope and diversity of available data. Existing datasets predominantly focus on written language and limited participant populations, which restrict ecological validity and generalizability. Expanding multimodal and multi-population linguistic neurodata therefore constitutes a key future priority.
Spoken language corpora aligned with neural recordings are especially important, as speech represents the primary mode of natural language use. Incorporating acoustic, articulatory, and prosodic information alongside neural signals can reveal encoding principles that are not captured by text-based paradigms alone. More broadly, integrating multimodal inputs such as vision and action can better reflect the conditions under which language is acquired and used in natural environments.
In parallel, large-scale brain imaging databases spanning diverse populations, languages, and developmental stages are needed. Such resources would enable systematic investigation of individual differences, cross-linguistic variation, and neurodiversity in language processing, providing a stronger empirical foundation for building and evaluating brain-aligned language models.
5.2.4 Establishing Multi-dimensional Evaluation Frameworks for Brain–Language Alignment
Current evaluation of language–brain alignment relies heavily on neural prediction accuracy, which provides only a partial view of cognitive and mechanistic validity. Establishing multi-dimensional evaluation frameworks is therefore essential for assessing genuine progress.
Future evaluation systems should incorporate metrics related to memory consolidation, including stability–plasticity trade-offs and long-term information retention, to assess whether models capture key properties of human learning. In addition, mechanistic consistency models can be developed to quantify whether internal model dynamics align with known neural principles, such as hierarchical processing, temporal integration, and causal information flow.
By jointly considering behavioral performance, neural predictability, cognitive plausibility, and mechanistic consistency, such multi-evaluation frameworks can provide a more rigorous and interpretable assessment of brain–language alignment, moving the field beyond surface-level correspondence toward deeper explanatory understanding.
6 Conclusion
Linguistics and neuroscience have traditionally advanced along largely independent trajectories, with the former emphasizing formal descriptions of language competence and the latter focusing on the biological mechanisms of language processing. Computational neuroscience offers an operational bridge by translating linguistic constructs into computable representations that can be quantitatively compared with neural signals. Historically, interdisciplinary progress has shifted from descriptive fitting toward mechanistic alignment. Early studies relied on simplified models with limited explanatory power, whereas recent advances in deep learning and large language models have enabled large-scale model–brain alignment by capturing both behavioral and neural patterns of language processing. However, such alignment remains largely correlational, as similarities may reflect shared sensitivity to statistical structure rather than shared computational mechanisms. Addressing this ambiguity requires alignment approaches constrained by both functional and structural principles, integrating causal analysis, model dissection, and biologically grounded constraints on neural architecture and dynamics. Looking forward, computational neuroscience has the potential to develop from an auxiliary methodology into a generative theoretical framework capable of producing testable neuro-computational accounts of language. Through the integration of linguistic theory, computational modeling, and neural evidence, it provides a necessary path toward biologically grounded explanations of human language.
Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund (Project ID: P0053210), The Hong Kong Polytechnic University Faculty Reserve Fund (Project ID: P0053738), an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428), The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097) and Research Grants Council Collaborative Research Fund (Ref: C5033-24G).